# RuleR: Improving LLM Controllability by Rule-based Data Recycling

### v1

@@ 3.2 Experimental Results
Furthermore, our method not only improves the constraint controllability but also keeps the general instruction-following ability of the original data. The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models. We hypothesize that the additional constraints might complicate the original instructions, thus leading to potentially improved instruction-following abilities.
@@ 3.3 Ablation Studies
Table 2: Ablations. The experiments are conducted on Mistral-7B on Alpaca-GPT4 data and the configuration notated with “(*)” represents our default setting. “IF Eval” and “Open LLM” represent the averaged scores.
In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset.
As shown in Table 2, “Single Temp.” represents utilizing only one rule-insturtcion template for each rule, while “Diverse Temp.” represents utilizing several different templates (with the same meaning) for each rule and those templates are randomly sampled during the augmentation. The superiority of “Diverse Temp.” showcases the necessity of providing diverse templates to avoid potential overfitting. “Max Rule = x” represents the setting in which at most x different rules can be sampled and utilized on each original data sample. The performances of models, whose Max Rule is set under 5, are promising, while the performance declines on all 3 evaluation metrics when the Max Rule is set as 5. These results indicate that too many rules utilized together in one sample might be harmful to LLM capability. “Aug Rate = x” represents there is a probability of x to apply our augmentation to each sample. It can be observed that increasing the augmentation rate is beneficial to the controllability of LLMs, while slightly harmful to the general instruction-following capability. However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
@@ 4 Conclusion
In this work, we proposed a novel method, Rule-based Data Recycling (RuleR) which utilizes the constraints derived from the existing data to enhance the dataset and incorporates these constraints into the training process, thereby improving the multi-constraint controllability of LLMs. This approach presents a promising direction for developing SFT-augmenting methods to improve LLM controllability, paving the way for more efficient advancements of LLM.
@@
Our method focuses on improving LLM controllability by rule-based data augmentation to avoid extra regeneration processes from humans or models. Though the alleviation of reliance on human/model editing is our main contribution, it inevitably limits the types of constraints that can be implemented on the original data sample. Currently, all the constraint rules are based on verifiable shallow syntactic characteristics like the occurrence and frequency of words or sentences, etc, lacking constraint rules that focus more on the semantic feature or content. Thus, our method shows great potential to be further enhanced if further involves small models, which will still remain efficient and avoid the regeneration process.
@@ Appendix B Evaluation Metrics
We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard
IFEval Zhou et al. (2023b) is the primary evaluation metric employed in our study due to its compatibility with our motivation. It focuses on evaluating how LLMs follow various additional constraints, such as specifying a word count or requiring the inclusion of certain keywords a specified number of times. To avoid the utilization of LLMs during evaluation, it proposes 25 distinct types of verifiable instructions. There are 541 prompts in total and each of them incorporates one or more of these verifiable instructions, ensuring the comprehensiveness of the evaluation. IFEval serves as a great inspiration for our method, and there exist semantical overlappings between their verifiable instructions and our rules. However, the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential data leakage. Moreover, IFEval only needs to verify the responses, while our method needs to modify the responses for the training, which pushes this process a step further. Consequently, this benchmark not only facilitates a comprehensive comparison but also provides valuable insights that align with our purpose.
Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts. This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al. (2024); Li et al. (2023b); Sottana et al. (2023). We utilize test instruction sets from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023b), comprising 218 and 80 diverse human-curated instructions, respectively. Following the framework by Li et al. (2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy. To mitigate positional bias Ko et al. (2020); Wang et al. (2023a), responses are presented in two distinct sequences. Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction. One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions. The specific evaluation prompt is presented in Figure 2.
@@ Appendix C Predefined Rules
Rule Type Rule Name Example Rule-Instruction Templates
@@
Table 3: The list of predefined constraint rules.

### v3

@@ 4.1 Main Results
Furthermore, our method not only improves the constraint controllability but also keeps the general instruction-following ability of the original data. The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models. We hypothesize that the additional constraints largely complicate the original instructions, as shown in Figure 3, thus forcing the LLMs to understand each constraint before generating responses, thus leading to potentially improved instruction-following abilities.
@@ 4.2 Ablation Studies
Table 2: Ablation Study. Average scores on “IF Eval” and “Open LLM” of Mistral-7B finetuned with ours-augmented Alpaca-GPT4 data. “(*)” represents default configuration.
Sub-Category Case Combination Content Json Keywords Language Length Punctuation Start End
Baseline (Strict) 22.47 16.92 75.47 63.06 44.79 58.06 31.47 12.12 58.21
Aug Rate = 0.1 78.65 63.08 45.28 59.24 33.13 74.19 32.17 30.30 79.10
Aug Rate = 0.3 76.40 58.46 45.28 45.86 30.06 58.06 35.66 22.73 71.64
Aug Rate = 0.5 70.79 69.23 49.06 45.22 39.88 77.42 28.67 13.64 61.19
Aug Rate = 0.7 74.16 61.54 43.40 47.77 33.13 51.61 30.77 77.27 79.10
Aug Rate = 0.9 67.42 75.38 47.17 47.13 34.97 61.29 31.47 66.67 82.09
Baseline (Loose) 24.72 20.00 75.47 66.88 48.47 64.52 37.06 15.15 58.21
Aug Rate = 0.1 79.78 69.23 45.28 61.15 37.42 74.19 36.36 40.91 80.60
Aug Rate = 0.3 77.53 66.15 45.28 47.13 36.81 64.52 39.16 25.76 76.12
Aug Rate = 0.5 77.53 70.77 49.06 45.86 42.94 80.65 32.17 15.15 65.67
Aug Rate = 0.7 77.53 70.77 43.40 48.41 35.58 58.06 35.66 84.85 80.60
Aug Rate = 0.9 70.79 80.00 47.17 48.41 40.49 67.74 39.16 69.70 85.07
Table 3: Sub-category performance on IFEval benchmark of Mistral-7B finetuned with RuleR-augmented Alpaca-GPT4 data. The top section represents the performance by the strict criterion while the bottom represents the loose.
In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset. The experiments aim to evaluate the impact of several factors on our method: template diversity, the maximum number of rules utilized in each data sample, and the probability of applying augmentation to each sample. The performance of each configuration is evaluated on three evaluation metrics, Instruction Following Eval Pair-wise winning scores, and Open LLM Leaderboard, to simultaneously inspect the constraint controllability and general instruction-following ability.
4.2.1 Effect of Templates
This section illustrates the effectiveness of utilizing diverse rule templates for the augmentation. As shown in Table 2, “Single Temp.” represents utilizing only one rule-instruction template for each rule, while “Diverse Temp.” represents utilizing several different templates, approximately $10$, with the same meaning for each rule and those templates are randomly sampled during the augmentation.
As indicated in the table, “Diverse Temp.” demonstrates superior performance in both IF Eval and Pair-wise winning scores, with slightly lower accuracy on the Open LLM leaderboard. This result suggests that “Diverse Temp.” enhances the model’s constraint controllability while enhancing its general instruction-following capabilities compared to “Single Temp.” On the contrary, when fixing the rule templates to one single template, potential overfitting to the template might occur and thus negatively influence the model performances.
4.2.2 Effect of Rule Numbers
This section illustrates the effects of implementing multiple random rules together in one single sample. “Max Rule = $x$” represents the setting in which at most $x$ different rules can be sampled and utilized on each original data sample. In the augmentation process, a random value will be sampled in the range of $[0,x]$ as the number of rules in this sample. The examples in Figure 2 showcase the scenario when only one rule is implemented and examples in Figure 3 showcase the scenario when multiple rules are implemented.
Compared to the baseline, nearly all settings show performance improvements across the three evaluation metrics. However, as shown in the table, the IF Eval score initially increases, reaching its peak when “Max Rule = $3$”, before declining. The Pair-wise score, on the other hand, consistently decreases from “Max Rule = $1$” to “Max Rule = $5$”. Although no clear trend is observed on the Open LLM leaderboard, Max Rule = 5 shows the worst performance. These results suggest that applying too many rules to a single sample may impair the LLM’s capability, even though sampling and applying multiple rules can be beneficial when done in moderation, which might be because the original instruction becomes so complex for LLMs to learn.
4.2.3 Effect of Augmentation Rate
This section illustrates the effect of the augmentation rate in our method. “Aug Rate = $x$” represents there is a probability of $x$ to apply our augmentation to each sample.
It is observed that as the augmentation rate increases, performance improves on the IF Eval and mostly improves on the Pair-wise evaluation. However, the results on the Open LLM leaderboard show limited improvement from the baseline, without a significant upward trend. This phenomenon indicates that increasing the augmentation rate primarily enhances the LLM’s constraint controllability, and it also has a positive impact on its general instruction-following capability. However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
4.3 Detailed Sub-Category Analysis
In this section, detailed comparisons between our models and baseline models are provided in Table 3 for further analysis of the effects of our method. It is worth noting that the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential overfitting to the templates.
Models trained with our recycled data outperform the baseline model consistently in “Case”, “Combination”, “Punctuation”, and “Start End” categories, which are partially well-recycled by our method. In the “Length” category, our models are only slightly better than the baseline model although our recycling method contains this kind of constraints. After further investigation, we find the performance in this category is mainly influenced by the original data length distributions. Since our method does not introduce more data samples, thus not able to improve the performance dramatically, but the better performance indeed provides the model with a better understanding of response length. Interestingly, the performance in the “Language” category also shows consistent improvement although we do not introduce any more data. Considering the consistency between this performance and the Pair-wise comparison performance, we hypothesize this improvement is caused by the better general instruction-following abilities provided by our method.
The performance in the “Content” category presents one of the limitations of our Rule-based Recycling method, without utilizing other models or human experts to rewrite the instruction and response, it’s hard for our method to modify the content of the existing response. The performances of our models in the “Json” and “Keywords” categories are merely slightly lower, which is mainly affected by the diversity of original training datasets.
Comparing the performance changes across augmentation rates, LLMs obtain better performances when the augmentation rate is higher, except for “Case”, indicating the easiness of case-related constraints for LLMs to understand and learn.
@@ 5 Conclusion
In this work, we proposed Rule-based Data Recycling (RuleR), which modified the original instructions and responses from an existing dataset by rule-defined constraints. The “recycled” data aims to enhance the LLMs’ capability to generate outputs fulfilling the constraints specified in the input, thereby improving the controllability of LLMs. RuleR took the first step of exploring rule-based data recycling. As a data generation approach, it is much cheaper than existing methods generating new data from scratch since RuleR does not rely on any human supervision or additional models. Our RuleR can serve as a plug-and-play and easy-to-use method that converts any existing instruction-tuning datasets to new datasets for gaining better controllability. Hence, our method is more accessible and widely applicable. RuleR provides a promising and sustainable strategy to continuously improve LLMs. Moreover, our method has great potential to be extended to other settings. For example, with the help of additional models, we can apply content-level or semantic-level rules to produce more diverse training data, as discussed in the next section.
@@
Our method focuses on improving LLM controllability by rule-based editing of existing data, thereby avoiding the extra cost of data generation by humans or expert models. Though it saves the cost of human/model editing, the rules inevitably limit the types of constraints that can be applied to modify the original data. In the presented RuleR, all the constraints and rules are based on verifiable shallow syntactic characteristics such as the occurrence and frequency of words or sentences while lacking constraints and controllability on the semantic feature or content. This implies a potential of RuleR to be further enhanced by modifying the semantic content with a smaller model, which retains a comparable efficiency of rule-based editing.
@@ Appendix B Evaluation Metrics
The prompt for pair-wise comparison is shown in Figure 4.
@@ Appendix C Predefined Rules
Rule Type Rule Name Example of an Instruction Template for the Rule
@@
Table 4: The list of predefined constraint rules. Each rule contains (1) a set of constraint templates that serve as additional rule-instructions, on constraints the LLM should follow, and (2) specified methods that alternately edit the instruction and response to reach an alignment between them.